7. Assume you want to construct a portfolio from the following two securities. What is the expected return on your portfolio if you invest $800,000 in Security 1 and $200,000 in Security 2? Security Expected Return Beta 1 14.0% 1.12 2 10.5% 0.94

A. 13.8% B. 13.6% C. 13.3% D. 12.9% E. 12.3%
Answer #1

Total investment=(800,000+200,000)=1,000,000

Hence portfolio return=Respective return*Respective weight

=(800,000/1,000,000*14)+(200,000/1,000,000*10.5)

=13.3%